OpenStudy (anonymous):

If E and F are independent events, find P(F) if P(E) = 0.5 and P ( E u F) = 0.8. Round your answer to 3 decimal places. P(F) =

OpenStudy (anonymous):

\[P(E \cup F)=P(E)+P(F) - P(E and F)\]Here P(E and F) is zero since the events are independent. So you have\[P(E \cup F)=P(E)+P(F) \rightarrow P(F)=P(E \cup F)-P(E)=0.8-0.5=0.3\]

OpenStudy (anonymous):

To three decimal places, 0.300

OpenStudy (anonymous):

You have to remember the first formula to get anywhere.

OpenStudy (anonymous):

when i put that in as an answer it says it is wrong

OpenStudy (anonymous):

Hmm, I'm pretty sure that's the answer. Is this an online test or something?

OpenStudy (anonymous):

no my probability class works off of a thing called wiley plus it gives you questions and i have four tries to get it

OpenStudy (anonymous):

Hang on...

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Sorry...only had a few hours sleep and distracted...what we should have done is realized:\[P(E and F)=P(E)P(F)\]which isn't zero.

OpenStudy (anonymous):

No problem lol as long as you can help me some

OpenStudy (anonymous):

Technically,\[P(A|B)=\frac{P(A and B)}{P(B)} \rightarrow = P(A|B)P(B)=P(A and B)\]

OpenStudy (anonymous):

Now, P(A|B) = P(A) since it's independent of B, so P(E and F) = P(E)P(F) here.

OpenStudy (anonymous):

Then\[P(E \cup F) = P(E)+P(F)+P(E)P(F)=P(E)+P(F)(1+P(E))\]

OpenStudy (anonymous):

\[\frac{P(E \cup F)-P(E)}{1+P(E)}=P(F)\]

OpenStudy (anonymous):

Are these formulas coming through?

OpenStudy (anonymous):

somewhat I am trying to see the numbers going with it but im just confusing myself

OpenStudy (anonymous):

P(F)=[P(E or F)-P(E)]/[1-P(E)]

OpenStudy (anonymous):

so P(F)=[0.8 - 0.5]/[1-0.5] = 0.3/0.5 = 0.600

OpenStudy (anonymous):

How many goes on your Wiley thing do you have left?

OpenStudy (anonymous):

I had one more and that answer was correct so now i have to just figure out i think more problems and i will be good thanks so much for your help

OpenStudy (anonymous):

good! i'm relieved...

OpenStudy (anonymous):

because I made another mistake typing out, though on my notepad it's right...good luck

OpenStudy (anonymous):

fan me if you like :) need the points.

OpenStudy (anonymous):

okay no problem take a look at some of the other problems i have up if you have the time

OpenStudy (anonymous):

If I can, I will :)